OpenStudy (anonymous):

Please help! If the numerator of a fraction is increased by 3, the fraction becomes 3/4. If the denominator is decreased by 7, the fraction becomes 1. Determine the original fraction. Which of the following equations represents "If the numerator of a fraction is increased by 3, the fraction becomes 3/4"? (Hint: cross products) 3n + 9 = 4d 4n + 3 = 3d 4n + 12 = 3d @mathmale

OpenStudy (mathmale):

this problem statements translates into TWO different equations which must be solved simultaneously. Let the fraction be n/d (numerator over denominator). "Increase the numerator by 3 and find that the result equals 3/4" See whether you can write one equation that reflects this. share your equation.

OpenStudy (anonymous):

Would it be n+ 3 = 3/4

OpenStudy (anonymous):

@mathmale

OpenStudy (anonymous):

@rubyyy

OpenStudy (anonymous):

r u still there?

OpenStudy (anonymous):

xx+7 if the numerator is increased by 3, then x+3x+7=34 by cross multiplication we get, 4 ( x +3) = 3 (x+7) 4x+12=3x+21 4x-3x=21-12 x=9 So the required fraction is 9 / 16

OpenStudy (anonymous):

okay I think I get it. Thanks so much.

OpenStudy (anonymous):

No prob :P